You purchased a stock for $50 per share and sold it one year later for $60. Over that time period, it also paid dividends totaling $2 per share. What were your dollar return and your percentage return on the investment?

  1. $10 and 20% because you don’t include dividends in returns.
  2. $12 and 24% because you do include dividends in returns.
  3. $12 and 20% because you include dividends in dollar returns, but not percentage returns.
  4. It depends on whether arithmetic or geometric mean is used.
Answer #1

Dollar return = Selling price+dividend-purchase price = 60+2-50 = 12

Percentage return = Dollar return/purchase price = 12/50=24%